# Grading Massive Open Online Courses Using Large Language Models

### v1

@@ Abstract
Massive open online courses (MOOCs) offer free education globally to anyone with a computer and internet access. Despite this democratization of learning, the massive enrollment in these courses makes it impractical for one instructor to assess every student’s writing assignment. As a result, peer grading, often guided by a straightforward rubric, is the method of choice. While convenient, peer grading often falls short in terms of reliability and validity. In this study, we explore the feasibility of using large language models (LLMs) to replace peer grading in MOOCs. Specifically, we use two LLMs, GPT-4 and GPT-3.5, across three MOOCs: Introductory Astronomy, Astrobiology, and the History and Philosophy of Astronomy. To instruct LLMs, we use three different prompts based on the zero-shot chain-of-thought (ZCoT) prompting technique: (1) ZCoT with instructor-provided correct answers, (2) ZCoT with both instructor-provided correct answers and rubrics, and (3) ZCoT with instructor-provided correct answers and LLM-generated rubrics. Tested on 18 settings, our results show that ZCoT, when augmented with instructor-provided correct answers and rubrics, produces grades that are more aligned with those assigned by instructors compared to peer grading. Finally, our findings indicate a promising potential for automated grading systems in MOOCs, especially in subjects with well-defined rubrics, to improve the learning experience for millions of online learners worldwide.
@@
Shahriar Golchin, Nikhil Garuda, Christopher Impey, Matthew Wenger Department of Computer Science, University of Arizona, AZ, USA Department of Astronomy, University of Arizona, AZ, USA golchin@arizona.edu
@@ 1 Introduction
In this study, we investigate the potential of LLMs in handling the task of grading assignments in MOOCs and the viability of replacing the current peer grading system with LLMs on this platform, aiming to establish more personalized and automated grading feedback to enhance the online learning experience for millions of students globally. Specifically, we employ three distinct prompting strategies to assess the grading performance of two performant LLMs—GPT-4 (OpenAI, 2023) and GPT-3.5 (Ouyang et al., 2022)—in three MOOC subjects: Astrobiology (Impey, 2024a), Introductory Astronomy (Impey, 2024b), and the History and Philosophy of Astronomy (Impey, 2024c). Our methodology is primarily grounded on utilizing the zero-shot chain-of-thought (ZCoT) prompting technique Kojima et al. (2022) by asking the LLMs to grade assignments “step by step.” However, we tweak this prompt by providing more in-context information about the given task. In particular, we augment this prompt in a threefold manner: (1) ZCoT with instructor-provided correct answers, (2) ZCoT integrated with both instructor-provided correct answers and rubrics, and (3) ZCoT coupled with instructor-provided correct answers and LLM-generated rubrics based on the correct answers. Our findings suggest that GPT-4 using ZCoT with instructor-provided answers and rubrics outperforms peer grading in terms of score alignment with those given by instructors. However, grading courses that require imaginative or speculative thinking, such as History and Philosophy of Astronomy, is found to be a challenge for both LLMs and peer grading. Despite this, GPT-4 still excels over peer grading by producing scores more closely aligned with those assigned by the instructors in such scenarios.
@@
(1) For the first time, we study the potential of replacing the peer grading system with the use of LLMs in MOOCs, aiming for automated and more accurate grading feedback to improve the learning experience for online learners globally.
(2) Using a rigorous evaluation across 18 settings, our results reveal that when integrating ZCoT with instructor-provided correct answers and rubrics, along with GPT-4, it outperforms all other settings, including peer grading, in generating grades more aligned with those assigned by instructors. However, this alignment is more pronounced in courses that require less creative and imaginative thinking.
@@ 2 Related Work
The existing foundation for this research stems from preceding studies that explored peer grading in MOOCs (Formanek et al., 2017; Gamage et al., 2021). While they showed a connection between peer review participation and elements such as students’ engagement and course completion rates, they also reported issues related to inconsistency in grades and problems concerning reliability and validity within this process.
Regarding the exploration of LLMs in educational settings, the primary focus has been on their utilization by instructors and students to generate and respond to educational content (Kasneci et al., 2023; Alseddiqi et al., 2023). One such exemplary application is at Khan Academy where GPT-4 is employed as a learning assistance tool, aiding in student learning via content-specific question-answering (OpenAI, 2024).
Despite its relevance, there is a noticeable gap in research that directly parallels our study. Several studies (Morris et al., 2023; Funayama et al., 2022; Riordan et al., 2017) employed the classic encoder-only component of Transformer networks (Devlin et al., 2018; Vaswani et al., 2017), examining the grading of students’ assignments within short question-answering contexts. Despite their value, these works are limited to short question-answering settings and necessitate the training of individual, subject-specific models, all the while involving human interaction. The research bearing the closest resemblance to our current endeavor is (Pinto et al., 2023), wherein GPT-3.5 (integrated within the ChatGPT framework) is used in a corporate context, providing feedback on employee performance metrics.
Our work, while centers around grading students’ assignments, is substantially different from others in the sense that we are the first to use LLMs to study the possibility of replacing the peer grading system in MOOCs, aiming to mitigate the necessity for human involvement in the grading process while providing reliable and customized feedback.
@@ 3.1 Prompts
To guide LLMs in grading assignments, we use zero-shot chain-of-thought (ZCoT) (Kojima et al., 2022) as our base prompting technique. There are two main reasons behind this choice. First, we performed a pilot study and found that ZCoT yields scores more closely aligned with those given by instructors compared to vanilla zero-shot prompting. Second, ZCoT enables us to observe the reasoning behind each generated score, as the LLM is required to verbalize its thought process “step by step.” This allows us to thoroughly analyze the reliability and validity of the grading steps provided as feedback to students and proactively mitigate potential hallucinations during the grading process.
@@
Instruction: You are a fair and knowledgeable instructor whose task is to evaluate the student’s assignments in accordance with the correct answers to each of the questions that are presented in the section that follows. Make sure you speak your thoughts aloud so that the students can understand the rationale for any points deducted. Let’s grade assignments step by step. - - - Questions/Answers: Questions: What are the advantages of large telescopes? Provide at least one. Why do astronomers want telescopes in space when putting them there is expensive? What are some examples of wavelength regions beyond the spectrum of visible light where astronomers can learn about the universe? Provide at least two. Answers: Large telescopes collect more light and so permit fainter and more distant objects to be seen. Also, large telescopes also generally have higher angular resolution, although realizing this depends on being able to correct for blurring by the Earth’s atmosphere. Telescopes in space are immune from the blurring effect of the Earth’s atmosphere, which reduces the depth of observation and limits the ability to resolve or separate object close on the sky. For some wavelength regimes beyond the visible spectrum, astronomy must be done from space because those wavelengths do not penetrate the Earth’s atmosphere. Telescopes in space are also immune from light pollution on Earth. Wavelength ranges include: radio, microwave, infrared, ultraviolet, X-ray, and gamma ray. Grade: 9/9 - - - Student’s Answers: Astronomers can learn about the universe by investigating certain wavelengths such as cosmic and gamma rays, X-rays, infrared, and radio waves. For example, the Cosmic Background Radiation (CMB) can tell astronomers something about the recombination epoch. It is very important for observational cosmology. X-rays are produced in the universe when matter is heated to millions of degrees. These are therefore used to detect areas with high magnetic fields and extreme gravity plus areas where huge cosmic explosions occur. Likewise, gamma ray bursts (GRBs) occur when a massive star collapses to form a neutron star or black hole. […] {The answer is partially cut due to space constraints.} Grade: - - - 7/9 The student’s answer is quite comprehensive and shows a good understanding of the subject matter. However, the student did not directly answer the first question about the advantages of large telescopes. While they did mention that “The light-gathering power and the resolution of a telescope increase with increasing aperture”, they did not explicitly state that these are advantages of large telescopes. The student did a good job explaining why astronomers want telescopes in space, mentioning that they “give ultra-sharp optical images and to detect long and short wavelengths that cannot penetrate the Earth’s atmosphere.” The student also did not provide specific examples of wavelength regions beyond the spectrum of visible light where astronomers can learn about the universe, which was asked in the third question. They mentioned the “peeling back of the electromagnetic spectrum, revealing for the first time the invisible universe” but did not provide specific examples as requested. Therefore, 2 points are deducted for not directly answering the first question and not providing specific examples for the third question.
@@
Instruction: You are a fair and knowledgeable instructor whose task is to evaluate the student’s assignments in accordance with the correct answers and provided rubric to each of the questions that are presented in the section that follows. Make sure that the score deduction follows the provided rubric. Let’s grade assignments step by step and explain the reasons behind the point deduction given the rubric. - - - Rubrics: Does the writer answer the question correctly and express their thoughts clearly using supporting facts and relevant examples? 3 points: The writer answers the question correctly and the writer expresses their thoughts clearly, and uses supporting facts and relevant examples 2 points: The writer answers the question correctly, but does not explain it well. 0 points: The writer does not answer the question correctly Does the writer answer the question correctly and express their thoughts clearly using supporting facts and relevant examples? 3 points: The writer answers the question correctly and the writer expresses their thoughts clearly, and uses supporting facts and relevant examples 2 points: The writer answers the question correctly, but does not explain it well. 0 points: The writer does not answer the question correctly Does the writer answer the question correctly and the writer expresses their thoughts clearly using supporting facts and relevant examples? 3 points: The writer answers the question correctly and the writer expresses their thoughts clearly, and uses supporting facts and relevant examples 2 points: The writer answers the question correctly, but does not explain it well. The write only includes one wavelength instead of two. 0 points: The writer does not answer the question correctly Questions/Answers: {As Figure 1.} Grade: 9/9 - - - Student’s Answers: {As Figure 1.} Grade: - - - 6/9 The student does not directly answer the first question about the advantages of large telescopes. While they mention that telescopes improve on the light-gathering power of the eye and allow astronomers to resolve finer details, they do not specifically state these as advantages of large telescopes. This is worth 2 points instead of 3. The student answers the second question correctly and explains well why astronomers want telescopes in space. They mention that telescopes in space can give ultra-sharp optical images and detect long and short wavelengths that cannot penetrate the Earth’s atmosphere. This is worth 3 points. The student does not directly answer the third question about examples of wavelength regions beyond the spectrum of visible light. While they mention the peeling back of the electromagnetic spectrum, they do not provide specific examples of wavelength regions. This is worth 1 point instead of 3.
@@
Instruction: Your task is to design a rubric that addresses the following questions/answers. The followings are the homework assignments for the Astrobiology course. This course is designed for undergraduate students majoring in Astronomy. The rubric must be unbiased and adaptable, capable of fairly evaluating any kind of student writing assignment. The scoring breakdown for each question should be as follows: the score for Question 1 is 10, the score for Question 2 is 10, and the score for Question 3 is 10. Your thoughtful consideration in creating this rubric will ensure all students’ work is evaluated equitably and consistently. Make sure that the rubric will provide points that are whole numbered. Each rubric should be dedicated to each question separately. - - - Questions/Answers: Question 1: {As Figure 1.} Answer 1: {As Figure 1.} Full Score: 10/10 Question 2: Discuss how habitable zone range and spectral type are related. […] {The question is partially cut due to space constraints.} Answer 2: The habitable zone is defined as the range of distances from a star within which liquid water can exist on a planet. […] {The answer is partially cut due to space constraints.} Full Score: 10/10 Question 3: Discuss why/whether it is plausible that exobiology exists at all, in terms of exoplanet type, spectral type and orbital distance. […] {The question is partially cut due to space constraints.} Answer 3: The star around HD-127688 has a K spectral type. […] {The answer is partially cut due to space constraints.} Full Score: 10/10 - - - Rubric for Question 1: - Identification of detection methods (2 points): The student correctly identifies the detection methods used for each exoplanet. - Explanation of detection methods (2 points): The student provides a clear and accurate explanation of how each detection method works. - Identification of physical characteristics (2 points): The student correctly identifies the physical characteristics that can be learned from each set of data. - Explanation of physical characteristics (2 points): The student provides a clear and accurate explanation of why these physical characteristics can be learned from the data. - Identification of Earth-like exoplanet (2 points): The student correctly identifies one exoplanet as Earth-like. […] {The answer is partially cut due to space constraints.}
@@ 3.2 Evaluation of LLM-Assigned Grades
As previously noted, MOOCs cater to a large number of participants for each course, and our study aims to evaluate the effectiveness of LLMs in grading assignments for a substantial student population. However, due to the budget limitations of our project when using proprietary LLMs and the need for careful analysis of the grades assigned by LLMs to individual questions, we conduct the grading process on a manageable subset of the entire collection of students’ assignments (see Section 4: Data for further details). On the other hand, it remains crucial to analyze the performance of LLMs in grading a significant volume of assignments, representative of the regular situations in MOOCs. To address this need, we employ the bootstrap resampling Efron (1979); Efron and Tibshirani (1993).
Qualitatively, we compare LLM-assigned grades with those given by instructors (ground truth). Specifically, we examine the absolute difference in average grades per question for each course, considering both lenient and strict grading. The hypothesis we investigate here is that the difference between LLM-assigned and instructor-assigned grades should not be statistically significant, as determined through bootstrap resampling. In fact, this implies that the average grades given by both LLM and the instructor could be similar enough to be considered interchangeable. In a different regime, we also use bootstrap resampling to evaluate average scores for potential bias in LLM grading. This helps us understand how the overall average scores from LLMs compare to those from instructors, providing further insight into the reliability and consistency of LLM grading.

### v2

@@ Abstract
Massive open online courses (MOOCs) offer free education globally. Despite this democratization of learning, the massive enrollment in these courses makes it impractical for an instructor to assess every student’s writing assignment. As a result, peer grading, often guided by a straightforward rubric, is the method of choice. While convenient, peer grading often falls short in terms of reliability and validity. In this study, we explore the feasibility of using large language models (LLMs) to replace peer grading in MOOCs. To this end, we adapt the zero-shot chain-of-thought (ZCoT) prompting technique to automate the feedback process once the LLM assigns a score to an assignment. Specifically, to instruct LLMs for grading, we use three distinct prompts based on ZCoT: (1) ZCoT with instructor-provided correct answers, (2) ZCoT with both instructor-provided correct answers and rubrics, and (3) ZCoT with instructor-provided correct answers and LLM-generated rubrics. We tested these prompts in 18 different scenarios using two LLMs—GPT-4 and GPT-3.5—across three MOOCs: Introductory Astronomy, Astrobiology, and the History and Philosophy of Astronomy. Our results show that ZCoT, when augmented with instructor-provided correct answers and rubrics, produces grades that are more aligned with those assigned by instructors compared to peer grading. Finally, our findings indicate a promising potential for automated grading systems in MOOCs, especially in subjects with well-defined rubrics, to improve the learning experience for millions of online learners worldwide.
@@
Shahriar Golchinthanks: Corresponding author., Nikhil Garuda, Christopher Impey, Matthew Wenger Department of Computer Science, University of Arizona Department of Astronomy, University of Arizona golchin@arizona.edu
@@ 1 Introduction
In this study, we investigate the potential of LLMs in handling the task of grading assignments in MOOCs and the viability of replacing the current peer grading system with LLMs on this platform, aiming to establish more personalized and automated grading feedback to enhance the learning experience for online learners. Specifically, our methodology is grounded on employing the zero-shot chain-of-thought (ZCoT) prompting technique by asking the LLMs to grade assignments “step by step” Kojima et al. (2022). Although this technique was originally designed to enhance the reasoning abilities of LLMs, we adapt it to generate feedback to students by having the LLMs verbalize the reasons behind score deductions. We also further refine it by providing more in-context information about the given task. In particular, we augment this technique in three ways: (1) ZCoT with instructor-provided correct answers, (2) ZCoT integrated with both instructor-provided correct answers and rubrics, and (3) ZCoT coupled with instructor-provided correct answers and LLM-generated rubrics based on the correct answers. We use these prompting strategies to evaluate the grading performance of two performant LLMs—GPT-4 (OpenAI, 2023) and GPT-3.5 (Ouyang et al., 2022)—in three MOOC subjects: Astrobiology (Impey, 2024a), Introductory Astronomy (Impey, 2024b), and the History and Philosophy of Astronomy (Impey, 2024c). Our findings suggest that GPT-4 using ZCoT with instructor-provided answers and rubrics outperforms peer grading in terms of score alignment with those given by instructors. However, grading courses that require imaginative or speculative thinking, such as History and Philosophy of Astronomy, is found to be a challenge for both LLMs and peer grading. Despite this, GPT-4 still excels over peer grading by producing scores more closely aligned with those assigned by the instructors in such scenarios.
@@
(1) We study the potential of replacing the peer grading system with the use of LLMs and adapting the ZCoT prompting technique in MOOCs, aiming to provide automated grading feedback that is more accurate and reliable than peer grading.
(2) Using a rigorous evaluation across 18 settings, our results reveal that when integrating ZCoT with instructor-provided correct answers and rubrics, along with GPT-4, it outperforms all other settings, including peer grading. This approach results in grades that are more closely aligned with those provided by instructors and offers more useful feedback than peer grading. However, the score alignment is stronger in courses that require less creative and imaginative thinking.
@@ 2 Related Work
LLMs in Education. The use of LLMs in education has primarily focused on how instructors and students leverage them to generate and respond to educational content (Kasneci et al., 2023; Alseddiqi et al., 2023). A key example is the implementation of GPT-4 at Khan Academy, where it acts as a learning assistant, answering subject-specific questions (OpenAI, 2024). Recent studies expanded on this, exploring LLMs’ capabilities in providing feedback on open-ended questions and essays (Matelsky et al., 2023; Pinto et al., 2023; Meyer et al., 2024; Stahl et al., 2024), grading short-answer questions (Chang and Ginter, 2024), generating programming exercises (Del Carpio Gutierrez et al., 2024), and helping students with coding questions (Tripaldelli et al., 2024) or offering feedback on their code (Pankiewicz and Baker, 2023; Kiesler et al., 2023; Zhang et al., 2024; Gabbay and Cohen, 2024; Heickal and Lan, 2024; Jacobs and Jaschke, 2024). Before LLMs became widely used, research in this area relied on the encoder-only component of Transformer networks (Devlin et al., 2018; Vaswani et al., 2017) to evaluate assignments in short question-answering contexts (Morris et al., 2023; Funayama et al., 2022; Riordan et al., 2017).
LLMs as Judges. The most common use of LLMs is to approximate human judgment. Beyond grading, LLMs are used in various tasks requiring evaluative decisions. Among many examples, they were used to assess the quality of generated outputs in generation tasks such as summarization (Liu et al., 2023), evaluate translation quality (Kocmi and Federmann, 2023), and detect factual inconsistencies (Luo et al., 2023). LLMs were also applied to assess how well other LLMs follow instructions (Zeng et al., 2023), gauge the similarity between reference and LLM-generated texts (Golchin and Surdeanu, 2023b), and evaluate code (Zhuo, 2023). Building on this, fine-tuned LLMs as judges were proposed to enhance the performance of general-purpose LLMs (Zhu et al., 2023).
This research builds on earlier studies of peer grading in MOOCs, which revealed challenges in achieving consistency, reliability, and validity in peer-assigned grades (Formanek et al., 2017; Gamage et al., 2021). Our study explores the potential of using LLMs to replace peer grading in MOOCs, aiming to improve and automate the feedback and grading process.
@@ 3.1 Prompts
To guide LLMs in grading assignments, we use zero-shot chain-of-thought (ZCoT) (Kojima et al., 2022) as our base prompting technique. There are two main reasons behind this choice. First, we performed a pilot study and found that ZCoT yields scores more closely aligned with those given by instructors compared to vanilla zero-shot prompting. Second, ZCoT enables us to observe the reasoning behind each generated score, as the LLM is required to verbalize its reasoning process “step by step.” This allows us to analyze the reliability and validity of the grading steps provided as feedback to students and proactively mitigate potential hallucinations during the grading process. Note that our use of ZCoT differs from its conventional application. Typically, ZCoT involves reasoning before making predictions. However, in our approach, the score (prediction) is generated first, followed by reasoning as feedback on any score deductions. Thus, we use ZCoT not to enhance performance, but to automatically generate feedback.
@@
Instruction: You are a fair and knowledgeable instructor whose task is to evaluate the student’s assignments in accordance with the correct answers to each of the questions that are presented in the section that follows. Make sure you speak your thoughts aloud so that the students can understand the rationale for any points deducted. Let’s grade assignments step by step. – – – Questions/Answers: Questions: What are the advantages of large telescopes? Provide at least one. Why do astronomers want telescopes in space when putting them there is expensive? What are some examples of wavelength regions beyond the spectrum of visible light where astronomers can learn about the universe? Provide at least two. Answers: Large telescopes collect more light and so permit fainter and more distant objects to be seen. Also, large telescopes also generally have higher angular resolution, although realizing this depends on being able to correct for blurring by the Earth’s atmosphere. Telescopes in space are immune from the blurring effect of the Earth’s atmosphere, which reduces the depth of observation and limits the ability to resolve or separate object close on the sky. For some wavelength regimes beyond the visible spectrum, astronomy must be done from space because those wavelengths do not penetrate the Earth’s atmosphere. Telescopes in space are also immune from light pollution on Earth. Wavelength ranges include: radio, microwave, infrared, ultraviolet, X-ray, and gamma ray. Grade: 9/9 – – – Student’s Answers: Astronomers can learn about the universe by investigating certain wavelengths such as cosmic and gamma rays, X-rays, infrared, and radio waves. For example, the Cosmic Background Radiation (CMB) can tell astronomers something about the recombination epoch. It is very important for observational cosmology. X-rays are produced in the universe when matter is heated to millions of degrees. These are therefore used to detect areas with high magnetic fields and extreme gravity plus areas where huge cosmic explosions occur. Likewise, gamma ray bursts (GRBs) occur when a massive star collapses to form a neutron star or black hole. […] {The answer is partially cut due to space constraints.} Grade: – – – 7/9 The student’s answer is quite comprehensive and shows a good understanding of the subject matter. However, the student did not directly answer the first question about the advantages of large telescopes. While they did mention that “The light-gathering power and the resolution of a telescope increase with increasing aperture”, they did not explicitly state that these are advantages of large telescopes. The student did a good job explaining why astronomers want telescopes in space, mentioning that they “give ultra-sharp optical images and to detect long and short wavelengths that cannot penetrate the Earth’s atmosphere.” The student also did not provide specific examples of wavelength regions beyond the spectrum of visible light where astronomers can learn about the universe, which was asked in the third question. They mentioned the “peeling back of the electromagnetic spectrum, revealing for the first time the invisible universe” but did not provide specific examples as requested. Therefore, 2 points are deducted for not directly answering the first question and not providing specific examples for the third question.
@@
Instruction: You are a fair and knowledgeable instructor whose task is to evaluate the student’s assignments in accordance with the correct answers and provided rubric to each of the questions that are presented in the section that follows. Make sure that the score deduction follows the provided rubric. Let’s grade assignments step by step and explain the reasons behind the point deduction given the rubric. – – – Rubrics: Does the writer answer the question correctly and express their thoughts clearly using supporting facts and relevant examples? 3 points: The writer answers the question correctly and the writer expresses their thoughts clearly, and uses supporting facts and relevant examples 2 points: The writer answers the question correctly, but does not explain it well. 0 points: The writer does not answer the question correctly Does the writer answer the question correctly and express their thoughts clearly using supporting facts and relevant examples? 3 points: The writer answers the question correctly and the writer expresses their thoughts clearly, and uses supporting facts and relevant examples 2 points: The writer answers the question correctly, but does not explain it well. 0 points: The writer does not answer the question correctly Does the writer answer the question correctly and the writer expresses their thoughts clearly using supporting facts and relevant examples? 3 points: The writer answers the question correctly and the writer expresses their thoughts clearly, and uses supporting facts and relevant examples 2 points: The writer answers the question correctly, but does not explain it well. The write only includes one wavelength instead of two. 0 points: The writer does not answer the question correctly Questions/Answers: {As Figure 1.} Grade: 9/9 – – – Student’s Answers: {As Figure 1.} Grade: – – – 6/9 The student does not directly answer the first question about the advantages of large telescopes. While they mention that telescopes improve on the light-gathering power of the eye and allow astronomers to resolve finer details, they do not specifically state these as advantages of large telescopes. This is worth 2 points instead of 3. The student answers the second question correctly and explains well why astronomers want telescopes in space. They mention that telescopes in space can give ultra-sharp optical images and detect long and short wavelengths that cannot penetrate the Earth’s atmosphere. This is worth 3 points. The student does not directly answer the third question about examples of wavelength regions beyond the spectrum of visible light. While they mention the peeling back of the electromagnetic spectrum, they do not provide specific examples of wavelength regions. This is worth 1 point instead of 3.
@@
Instruction: Your task is to design a rubric that addresses the following questions/answers. The followings are the homework assignments for the Astrobiology course. This course is designed for undergraduate students majoring in Astronomy. The rubric must be unbiased and adaptable, capable of fairly evaluating any kind of student writing assignment. The scoring breakdown for each question should be as follows: the score for Question 1 is 10, the score for Question 2 is 10, and the score for Question 3 is 10. Your thoughtful consideration in creating this rubric will ensure all students’ work is evaluated equitably and consistently. Make sure that the rubric will provide points that are whole numbered. Each rubric should be dedicated to each question separately. – – – Questions/Answers: Question 1: {As Figure 1.} Answer 1: {As Figure 1.} Full Score: 10/10 Question 2: Discuss how habitable zone range and spectral type are related. […] {The question is partially cut due to space constraints.} Answer 2: The habitable zone is defined as the range of distances from a star within which liquid water can exist on a planet. […] {The answer is partially cut due to space constraints.} Full Score: 10/10 Question 3: Discuss why/whether it is plausible that exobiology exists at all, in terms of exoplanet type, spectral type and orbital distance. […] {The question is partially cut due to space constraints.} Answer 3: The star around HD-127688 has a K spectral type. […] {The answer is partially cut due to space constraints.} Full Score: 10/10 – – – Rubric for Question 1: - Identification of detection methods (2 points): The student correctly identifies the detection methods used for each exoplanet. - Explanation of detection methods (2 points): The student provides a clear and accurate explanation of how each detection method works. - Identification of physical characteristics (2 points): The student correctly identifies the physical characteristics that can be learned from each set of data. - Explanation of physical characteristics (2 points): The student provides a clear and accurate explanation of why these physical characteristics can be learned from the data. - Identification of Earth-like exoplanet (2 points): The student correctly identifies one exoplanet as Earth-like. […] {The answer is partially cut due to space constraints.}
@@ 3.2 Evaluation of LLM-Assigned Grades
As noted, MOOCs cater to a large number of participants for each course, and our study aims to evaluate the effectiveness of LLMs in grading assignments for a substantial student population. However, to control the budget when using proprietary LLMs, and carefully analyze the grades assigned by LLMs to individual questions, we conduct the grading process on a manageable subset of the entire collection of students’ assignments (see Section 4 under Data for more details). On the other hand, it is crucial to analyze the performance of LLMs in grading a significant volume of assignments, representative of the regular situations in MOOCs. To address this, we use the bootstrap resampling Efron (1979); Efron and Tibshirani (1993).
Qualitatively, we compare LLM-assigned grades with those given by instructors (ground truth). Specifically, we examine the absolute difference in average grades per question for each course, considering both lenient and strict grading. The hypothesis we investigate here is that the difference between LLM-assigned and instructor-assigned grades should not be statistically significant, as determined through bootstrap resampling. In fact, this implies that the average grades given by both LLM and the instructor could be similar enough to be considered interchangeable. In a different analysis, we also use bootstrap resampling to evaluate average scores for potential bias in LLM grading. This helps us understand how the overall average scores from LLMs compare to those from instructors, providing insight into the reliability and consistency of LLM grading. In addition, we compute the mean absolute error (MAE) to measure the discrepancies between LLM- and instructor-assigned grades.